Determine the percentage composition of each element in Na2PO3
1
Expert's answer
2012-11-09T02:38:06-0500
M(Na2PO3) = 23*2 + 31 + 16*3 = 125 g/mol
w(Na) = Ar * n / M *100%
w(Na) = 2*23 / 125 * 100% = 36.8 %
w(P) = 31/125 * 100% = 24.8 %
w(O) = 16*3 /125 * 100% = 38.4 %